undeadknight26 (undeadknight26):

Derive the equation of the parabola with a focus at (-2, 4) and a directrix of y = 6. Put the equation in standard form. @Vickiesweet

undeadknight26 (undeadknight26):

I think its B.

undeadknight26 (undeadknight26):

f(x) = one fourthx2 - x + 4 f(x) = -one fourthx2 - x + 4 f(x) = one fourthx2 - x + 5 f(x) = -one fourthx2 - x + 5

OpenStudy (paxpolaris):

the vertex is at (-2,5) \[y=-{1\over4(1)}\left( x+2 \right)^2+5\]

OpenStudy (paxpolaris):

i think you are correct